any help on this question would be greatly appreciated! thanks! :)
An object is dropped, its speed continually increases, and therefore its height above the ground decreases at a faster and faster rate. the height h (in feet) of the object t seconds after it is dropped is modeled by the function h = -16t^2 + h0 where h0 is the objects initial height. niagara falls in ny is 167 feet high. how long does it take for the water to fall from the top to the bottoms of the falls?

Answer by stanbon(75887)

An object is dropped, its speed continually increases, and therefore its height above the ground decreases at a faster and faster rate. the height h (in feet) of the object t seconds after it is dropped is modeled by the function h = -16t^2 + h0 where h0 is the objects initial height. niagara falls in ny is 167 feet high. how long does it take for the water to fall from the top to the bottoms of the falls?
----------------
h(t) = -16t^2+ho
---
Height is zero when the water is at the bottom of the falls
Solve:
-16t^2+167 = 0
-16t^2 = -167
t^2 = 10.44
Positive solution:
t = sqrt(10.44) = 3.23 seconds
==================================
Cheers,
Stan H.
=========